OpenStudy (anonymous):

Find quotient please (12x^3-16x^2)/2x

OpenStudy (aum):

\[ \frac{12x^3-16x^2}{2x} = \frac{12x^3}{2x} - \frac{16x^2}{2x} = ~ ? \]

OpenStudy (anonymous):

so its -16x^2/2x ?